# L-FUSION: Laplacian Fetal Ultrasound Segmentation & Uncertainty Estimation

### v2

@@ Abstract
Accurate analysis of prenatal ultrasound (US) is essential for early detection of developmental anomalies. However, operator dependency and technical limitations (e.g. intrinsic artefacts and effects, setting errors) can complicate image interpretation and the assessment of diagnostic uncertainty. We present L-FUSION (Laplacian Fetal US Segmentation with Integrated FoundatiON models), a framework that integrates uncertainty quantification through unsupervised, normative learning and large-scale foundation models for robust segmentation of fetal structures in normal and pathological scans. We propose to utilise the aleatoric logit distributions of Stochastic Segmentation Networks and Laplace approximations with fast Hessian estimations to estimate epistemic uncertainty only from the segmentation head. This enables us to achieve reliable abnormality quantification for instant diagnostic feedback. Combined with an integrated Dropout component, L-FUSION enables reliable differentiation of lesions from normal fetal anatomy with enhanced uncertainty maps and segmentation counterfactuals in US imaging. It improves epistemic and aleatoric uncertainty interpretation and removes the need for manual disease-labelling. Evaluations across multiple datasets show that L-FUSION achieves superior segmentation accuracy and consistent uncertainty quantification, supporting on-site decision-making and offering a scalable solution for advancing fetal ultrasound analysis in clinical settings.
@@ 1 Introduction
Fetal ultrasound imaging is a crucial tool for monitoring fetal development and facilitating the early detection of potential abnormalities, enabling timely interventions that support optimal care. Detection rates for abnormalities have significantly improved, rising from $15-59~{}\%$ in the 2000s to $30-88~{}\%$ in recent years, depending on the type of lesion [6]. However, ultrasound images are noisy and contain shadows, variability and ambiguities in anatomy, creating challenges for clinicians to precisely identify subtle, early-stage, or rare abnormalities. Therefore, computerised support using machine learning models is expected to improve outcomes. Such models can mitigate operator dependence by providing consistent, objective assessments, and for selected developmental problems, their detection rates are often on par with human observations [5, 21]. However, these models, when designed to provide predictions without any confidence bounds, may lack the ability to convey their certainty to the operator. In a safety-critical domain such as healthcare, modelling uncertainty quantification can improve diagnostic accuracy and also assist clinicians in focusing their attention on critical regions by effectively flagging areas of high uncertainty. Two main constraints limit the application of those models: Firstly, methods for uncertainty estimation suffer from high computational costs [26], especially, with an increasing number of imaging dimensions. Laplacian uncertainty approximation emerges as a particularly powerful framework, enabling fine-grained distinctions between anatomical variations seen in healthy and pathological cases. Furthermore, by incorporating faster Hessian approximations the framework keeps computational requirements tractable and speeds up compute time. Secondly, the main challenge for effective probabilistic approaches and efficient uncertainty estimation has been their requirement for large datasets. Epistemic uncertainty captures gaps in model knowledge, while aleatoric uncertainty accounts for noise and inherent variability in the data [10]. Epistemic uncertainty can be decreased by increasing the amount of data but is required to capture situations not encountered in the training set. Especially, in safety-critical systems, epistemic uncertainty is used to detect situations which have never been seen by the model before [13]. However, data scarcity, along with challenges in robustness and generalization, continues to be critical issues in ultrasound imaging. Integrating pretrained large-scale foundation models [11] in our Laplacian framework can address these challenges by leveraging their ability to generalise across diverse datasets while minimising the need for extensive retraining. These models are believed to transfer knowledge from large-scale training data to specialised domains, reducing training costs while significantly improving diagnostic performance. The benefits are twofold: first, they improve detection rates for subtle and complex conditions, and second, they ensure adaptability to variable imaging environments and limited data scenarios. When combined with advanced uncertainty frameworks like Laplacian approximations, these systems are uniquely equipped to tackle the statistical and computational requirements of uncertainty estimation in challenging datasets.
Contributions: We propose L-FUSION, a novel approach that integrates Laplacian uncertainty approximation with embeddings from a foundation model to deliver accurate segmentations and precise uncertainty quantification for fetal ultrasound. The key contributions of this work are:
(1) We utilise a deterministic ultrasound foundation model and facilitate aleatoric and epistemic uncertainty quantification through Laplacian approximation of weight posteriors of a segmentation head with Gaussian distributions. Our method achieves superior segmentation performance in both binary and multi-class applications, while enhancing unsupervised out-of-distribution detection by incorporating additional uncertainty quantification and segmentation counterfactual through a dropout component.
(2) By restricting the computationally intensive Hessian calculation for the Laplace approximation to the segmentation head, we enhance our method by integrating a fast Hessian approximation with an efficient, versatile encoder. This results in a highly efficient and generalisable solution for universal ultrasound imaging tasks in clinical applications.
(3) We introduce a 3D extension of our method that combines a deterministic encoder with a Laplacian segmentation head for ultrasound clips. For comparison, we developed a 3D Laplacian segmentation model, as no universal ultrasound video foundation model currently exists for optimal benchmarking.
@@
Related Work: Despite its widespread clinical use, automated segmentation of ultrasound imaging lags behind other imaging modalities (e.g. lack of standardised performance measures tailored for US, smaller datasets, high inter- and intra-expert variability) [19]. However, there is a growing emphasis on better validation, including comparisons across methods and evaluation on larger, standardised datasets, to support broader clinical adoption [19]. Recent research in ultrasound segmentation has shifted towards more systematic use of priors, such as geometric, temporal, and imaging physics constraints, to improve segmentation accuracy [18, 23, 14, 8]. Foundation models, trained on large-scale natural image datasets, have been adapted for various medical imaging modalities, such as X-ray, computer tomography, magnetic resonance imaging, and ultrasound. These models learn modality-specific features, but their ability to generalise across modalities remains limited [25]. In ultrasound imaging, two notable foundation models have recently emerged to address challenges specific to this modality. The Multi-Organ FOundation (MOFO) [2] model seeks to overcome data scarcity and the limitations of single-organ models by segmenting images across multiple organs. By extracting organ-invariant representations and incorporating anatomical priors, MOFO enhances segmentation accuracy by exploiting relationships between anatomical structures from different regions. Evaluated on a multi-organ database, MOFO outperforms single-organ segmentation methods in various metrics. With a similar aim, the Universal ultrasound Foundation Model (USFM) [11] tackles generalisation issues across tasks and organs by training on a large-scale multi-organ, multi-centre, and multi-device database. USFM uses self-supervised pre-training with a spatial-frequency dual-masked image modelling method to extract robust features from low-quality images. Foundation models extract rich representations, but they remain indifferent to uncertainty. Bayesian modelling provides a probabilistic approach to segmentation, quantifying both epistemic and aleatoric uncertainty [15, 1, 3, 7]. Laplacian approximations can be employed within a Bayesian framework to estimate epistemic uncertainty by approximating the posterior distribution. However, the direct computation of the Hessian is infeasible, as it scales quadratically with the number of model parameters [4]. Recent work [24] combines the aleatoric logit distribution from Stochastic Segmentation Networks [17] with Laplace approximations for epistemic uncertainty quantification.
@@
Figure 1: Foundation model (Encoder) with Laplacian segmentation head $L(f)$ and a dropout unit. (a) $L(f)$ decoding the embeddings $f$ from the Foundation model $\bar{F}(x)$ and a Dropout U-Net $G(x)$. (b) $L(f)$ and Dropout Segmentation Head $G(f)$ decoding the embeddings $f=\bar{F}(x)$. For both approaches, the segmentations $S_{i}(f)$ are predicted by $L(f)$. We enhance the uncertainty maps $U$ with OOD Calibration by including the variance and total entropy of logit differences.
@@ 2 Method
As illustrated in Fig. 1 the L-FUSION framework integrates three key components: (1) USFM [11] as the encoder, (2) a Laplacian segmentation head and (3) a Dropout unit for combined segmentation and uncertainty quantification with enhanced uncertainty maps.
Laplacian Component Given an input fetal ultrasound image $x$, the pretrained foundation model processes $x$ to generate embeddings $f$. All weights remain deterministic and frozen (i.e., no further fine-tuning or randomisation), which significantly reduces the complexity and computational cost typically involved in training deep neural networks. The embeddings $f$ are then passed to the Laplacian segmentation head $L$, which produces a normal distribution of logits $\eta_{L}|f\sim\mathcal{N}(\mu(f,\theta_{1}),\varSigma(f,\theta_{2}))$. $\theta_{1}$ and $\theta_{1}$ parametrise the distribution of segmentation head networks. Next, we sample the segmentation mask $S_{i}$ from the Bernoulli distribution parametrised by the sigmoid of the respective logit $\eta_{L,i}$, for $i=1,...,n$ with $n$ samples per input. For aleatoric uncertainty, we Monte Carlo-sample $m$ segmentation networks $L_{j}(f)$ for $j=1,...,m$ from the Laplace approximation $q(\theta^{*})$ for an embedding $f=\bar{F}(x)$. Each sampled segmentation network predicts one logit distribution $p(\eta|f,\theta)$ for each embedding $f$. Finally, we get the mean prediction map $\hat{S}_{L}$ and the uncertainty maps $U_{L,k}$, for $k=1,...,K$ with K uncertainty measures, expressed as $\hat{S}_{L},U_{L,k}=L(\bar{F}(x))$, with the frozen encoder $\bar{F}$. $U_{L,k}^{E}$ reflects epistemic uncertainty by measuring the mutual information $\mathcal{I}(p(S,\theta^{*}|f,\mathcal{F}))$, with the entirety of embeddings $\mathcal{F}$, indicating the model’s confidence and highlighting areas with less reliable predictions. $U_{L,k}^{A}$ captures aleatoric uncertainty by calculating the expected entropy $\mathcal{H}(p(S|f,\theta^{*})$, which results from data noise such as poor imaging quality or natural variability in fetal anatomy. Given the low-rank matrix of the Hessian $P$, $p(\eta|f,\theta)=\mathcal{N}(\eta|\mu_{\theta},PP^{\top}+D)$ defines the low-rank parametrisation of the covariance matrix [17] and $p(\eta|f,\theta)=\mathcal{N}(\eta|\mu_{\theta},PP^{\top}+D)$ the fast Hessian approximations [24]. Both allow faster computation of the Laplace approximation.
@@
Let $x\in\mathcal{X}$ be an input image, and let $\bar{F}:\mathcal{X}\to\mathcal{F}$ be a deterministic, pretrained foundation model mapping $x$ to a feature embedding $f=\bar{F}(x)$. Consider a Laplacian segmentation head $L$ that models the posterior segmentation probability as $p(S|f,\theta)=\prod_{i}\text{Bernoulli}(s_{i}|\sigma(\eta_{L,i})),$ where the logit distribution follows the Laplacian approximation, $p(\eta|f,\theta)=\mathcal{N}(\eta|\mu_{\theta},PP^{\top}+D).$ If $\bar{F}$ preserves segmentation-relevant information, i.e., $I(S;x)\newline =I(S;f),$ then $f$ is a sufficient statistic for segmentation uncertainty estimation.
@@
Since $\bar{F}$ is deterministic, the segmentation posterior simplifies to $p(S|x)=p(S|f,\theta),$ which holds if $p(f|x)$ is a delta distribution centered at $\bar{F}(x)$, meaning $p(f|x)=\delta(f-\bar{F}(x)).$ By the Data Processing Inequality [20], $I(S;x)\geq I(S;f),$ and since $f$ is a deterministic function of $x$, we get $I(S;x)=I(S;f),$ ensuring that $f$ is a sufficient statistic for segmentation. For uncertainty estimation, the Laplacian approximation assumes a Gaussian weight posterior, $q(\theta^{*})=\mathcal{N}(\theta^{*}|\theta_{\text{MAP}},\mathbf{H}^{*-1}),$ which propagates uncertainty into the predictive variance: $\operatorname{Var}(S|f)=J\mathbf{H}^{*-1}J^{T}+\sigma^{2}(f)I.$ Here, $J\mathbf{H}^{*-1}J^{T}$ captures epistemic uncertainty, and $\sigma^{2}(f)$ captures aleatoric uncertainty. Since both uncertainties depend only on $f$, it follows that $f$ is sufficient for segmentation and uncertainty quantification.
Dropout Component Additionally, we use a Dropout unit, a Dropout U-Net $G(x)$ (Fig. 1 a) or a Dropout Segmentation Head $G(f)$ (Fig. 1 b) taking the embeddings from the Foundation Model encoder as inputs. From the Dropout unit, we take the logit distribution $\eta_{MC}$ for sampling logits and epistemic uncertainty measures. For uncertainty quantification, we propose an OOD-Calibration mechanism based on logit variance $U_{V}=Var(\eta_{q})$ (Intra-Model) or logit difference variance between $U_{V}=Var(\left(\frac{1}{q}\sum_{q=1}^{Q}\eta_{q}\right)$ for $Q$ models (Inter-Model), and scaled by OOD measure from the same model (Intra-Model) or other model(s) (Inter-Model) for improving OOD detection AUC. For OOD calibration, we generate the uncertainty maps by including the total entropy $\mathcal{H}(U_{V})$ through its Hadamard product with $U_{V}$ which scales and limits the uncertainty maps to only areas of total entropy larger than zero in $U_{V}$.

### v4

@@ Abstract
Accurate analysis of prenatal ultrasound (US) is essential for early detection of developmental anomalies. However, operator dependency and technical limitations (e.g. intrinsic artefacts and effects, setting errors) can complicate image interpretation and the assessment of diagnostic uncertainty. We present L-FUSION (Laplacian Fetal US Segmentation with Integrated FoundatiON models), a framework that integrates uncertainty quantification through unsupervised, normative learning and large-scale foundation models for robust segmentation of fetal structures in normal and pathological scans. We propose to utilise the aleatoric logit distributions of Stochastic Segmentation Networks and Laplace approximations with fast Hessian estimations to estimate epistemic uncertainty only from the segmentation head. This enables us to achieve reliable abnormality quantification for instant diagnostic feedback. Combined with an integrated Dropout component, L-FUSION enables reliable differentiation of lesions from normal fetal anatomy with enhanced uncertainty maps and segmentation counterfactuals in US imaging. It improves epistemic and aleatoric uncertainty interpretation and removes the need for manual disease-labelling. Evaluations across multiple datasets show that L-FUSION achieves superior segmentation accuracy and consistent uncertainty quantification, supporting on-site decision-making and offering a scalable solution for advancing fetal ultrasound analysis in clinical settings. Code is available at https://github.com/ividja/L-FUSION.
@@ 1 Introduction
Fetal ultrasound (US) is a key tool for monitoring development and detecting abnormalities early, with detection rates improving from $15\text{--}59\%$ in the 2000s to $30\text{--}88\%$ today, depending on the lesion type [6]. However, US imaging remains challenging due to noise, shadows, and anatomical ambiguity. Machine learning models can support clinicians by providing consistent and objective assessments, with detection performance often matching expert evaluation [5, 24]. Yet, standard deterministic models offer no confidence estimates, limiting trust and interpretability in safety-critical settings. Uncertainty quantification helps flag unreliable predictions and guide clinical attention. Current approaches, however, face two main challenges: (1) high computational cost, especially for high-dimensional imaging [29], and (2) the need for large, diverse datasets to capture epistemic uncertainty, critical for identifying unfamiliar or out-of-distribution (OOD) cases [12, 15]. Laplacian approximations offer a computationally efficient method for capturing both epistemic and aleatoric uncertainty, especially when paired with fast Hessian estimates. Meanwhile, pretrained US foundation models [13] address data scarcity by enabling transfer learning across domains. When integrated, these components provide accurate, generalisable segmentation with reliable uncertainty estimates, supporting robust and scalable diagnostic systems for real-world fetal US. Contributions. We propose L-FUSION, a novel framework for fetal ultrasound segmentation that integrates foundation model embeddings with two complementary uncertainty-aware heads. Our key contributions are: (1) We combine a deterministic ultrasound foundation encoder with a Laplacian head for calibrated uncertainty and a Dropout head for counterfactual segmentations, enabling joint aleatoric and epistemic uncertainty quantification. This improves segmentation accuracy and unsupervised OOD detection. (2) We reduce computational cost by applying Laplace approximation only to the segmentation head and using efficient Hessian approximations. The frozen encoder and lightweight heads ensure scalability across diverse clinical ultrasound tasks. (3) We introduce the first 3D Laplacian segmentation model for ultrasound clips, enabling spatiotemporal extension of our method. In the absence of 3D foundation models, we leverage the 2D encoder as a proxy.
@@
Related Work. Automated ultrasound segmentation lags behind other imaging modalities due to inconsistent metrics, smaller datasets, and high inter- and intra-expert variability [20]. However, growing emphasis on rigorous validation, cross-method comparisons, and larger standardised datasets aims to support clinical uptake [20]. Recent advances incorporate systematic priors, and geometric, temporal, and physics-based constraints to boost accuracy [19, 26, 16, 9]. Foundation models pretrained on large-scale natural images have been adapted for medical imaging, including ultrasound, though cross-modality generalisation remains limited [28]. The Multi-Organ FOundation (MOFO) model [2] enhances segmentation by learning organ-invariant representations and anatomical priors, outperforming single-organ approaches. Similarly, the Universal US Foundation Model (USFM) [13] improves generalisation across organs, devices, and centres via self-supervised spatial-frequency dual-masked image modelling on a large, diverse dataset. While foundation models offer rich features, they lack inherent uncertainty quantification. Bayesian methods provide probabilistic segmentation, estimating epistemic and aleatoric uncertainty [1, 3, 8]. Laplace approximations efficiently estimate epistemic uncertainty by approximating the posterior, avoiding costly Hessian computation [4]. Recent work combines aleatoric logit distributions from Stochastic Segmentation Networks [18] with Laplace approximations for comprehensive uncertainty quantification [27].
@@
Figure 1: Foundation model (Encoder) with Laplacian segmentation head $L(f)$ and a Dropout segmentation path. Here, $L(f)$ and Dropout Segmentation Head $G(f)$ decoding the embeddings $f=\bar{F}(x)$. The segmentations $S_{i}(f)$ are predicted by $L(f)$. We enhance the uncertainty measure maps $U_{L/D,k}$ with Uncertainty-Calibration as $U_{fused}$ by including the variance and total entropy of logits.
@@ 2 Method
We introduce L-FUSION, a probabilistic segmentation framework for fetal ultrasound that combines a frozen foundation model, e.g., USFM [13], with two complementary uncertainty-aware segmentation heads: a Laplacian-based branch and a Dropout-based branch [7]. By jointly modelling both epistemic and aleatoric uncertainty, L-FUSION produces more reliable predictions and highlights regions of ambiguity, particularly useful for clinical interpretation and decision-making. An overview of the framework is shown in Fig. 1.
Feature Encoding L-FUSION begins by passing an input $2D$ ultrasound image $x$ through a pretrained foundation encoder $\bar{F}$, e.g., USFM [13]. The encoder remains frozen during training to preserve the generalisable representations learned during pretraining and to reduce the computational burden associated with fine-tuning the encoder as well. The output feature embedding $f=\bar{F}(x)$ is shared by both segmentation branches. This design choice decouples representation learning from downstream uncertainty estimation, ensuring that all uncertainty arises from the segmentation heads rather than encoder variability.
@@
Let $x\in\mathcal{X}$ be an input image, and let $\bar{F}:\mathcal{X}\to\mathcal{F}$ be a deterministic, pretrained foundation model mapping $x$ to a feature embedding $f=\bar{F}(x)$. Consider a Laplacian segmentation head $L$ that models the posterior segmentation distribution as $p(S|f,\theta)=\prod_{i}\text{Bernoulli}(s_{i}|\sigma(\eta_{L,i})),$ where the latent logits $\eta$ follow a Laplace-approximated posterior: $p(\eta|f,\theta)=\mathcal{N}(\eta|\mu_{\theta}(f),PP^{\top}+D),$ with $\mu_{\theta}(f)$ the mean prediction, and $PP^{\top}+D$ the covariance matrix. If $\bar{F}$ preserves all segmentation-relevant information, i.e., using the mutual information $I(S;x)=I(S;f),$ then $f$ is a sufficient statistic for estimating segmentation and associated predictive uncertainty.
@@
Since $\bar{F}$ is deterministic, the distribution over $f$ given $x$ reduces to a delta function: $p(f|x)=\delta(f-\bar{F}(x)),$ so the segmentation posterior satisfies $p(S|x)=p(S|f,\theta).$ By the Data Processing Inequality [23], $I(S;x)\geq I(S;f).$ We assume $I(S;x)=I(S;f),$ i.e., that no information about $S$ is lost in passing from $x$ to $f$. This implies that $f$ is a sufficient statistic for $S$ with respect to $x$. Now, consider uncertainty estimation. The Laplace approximation assumes a Gaussian posterior over parameters: $q(\theta^{*})=\mathcal{N}(\theta^{*}|\theta_{\text{MAP}},\mathbf{H}^{*-1}),$ where $\mathbf{H}^{*}$ is the Hessian of the negative log-likelihood at the MAP estimate. Predictive uncertainty is then propagated via linearisation of the model output $\eta$ around $\theta_{MAP}$, yielding: $\operatorname{Var}(S|f)\approx J\mathbf{H}^{*-1}J^{\top}+\sigma^{2}(f)I,$ where $J$ is the Jacobian of $\mu_{\theta}(f)$ for $\theta$, evaluated at $\theta_{MAP}$, and $\sigma^{2}(f)$ is the predictive variance of the likelihood model (e.g., from $D$ in the Laplace covariance). The first term corresponds to epistemic uncertainty, and the second to aleatoric uncertainty. Since both terms depend only on $f$, and $f$ is sufficient for $S$, it follows that $f$ is also sufficient for segmentation uncertainty estimation.
Laplacian Segmentation Path. The first segmentation branch estimates uncertainty using a Laplace approximation centred on the MAP estimate of parameters $\theta_{\text{MAP}}$. Given embedding $f$ from the frozen encoder, the Laplacian segmentation head $L$ models logits as a Gaussian distribution, $\eta_{L}\mid f\sim\mathcal{N}(\mu(f,\theta_{1}),$ $\Sigma(f,\theta_{2}))$, where $\theta_{1}$ and $\theta_{2}$ govern the mean and covariance. Segmentation masks $S_{i}$ are sampled from Bernoulli distributions parameterised by the sigmoid of each logit $\eta_{L,i}$, for $i=1,\ldots,n$ samples per input. To capture aleatoric uncertainty from data noise and variability, Monte Carlo (MC) sampling [22] is performed over $m$ segmentation networks $L_{j}(f)$, $j=1,\ldots,m$, drawn from the Laplace approximation $q(\theta^{*})$ around the MAP estimate, using embedding $f=\bar{F}(x)$. Each sampled network outputs a logits distribution $p(\eta\mid f,\theta)$. Aggregating predictions yields the mean segmentation map $\hat{S}_{L}$ and uncertainty measures $U_{L,k}$, $k=1,\ldots,K$. $U_{L,k}^{E}$ quantifies epistemic uncertainty via mutual information $\mathcal{I}(p(S,\theta^{*}\mid f,\mathcal{F}))$ over embeddings $\mathcal{F}$, highlighting prediction unreliability. Conversely, $U_{L,k}^{A}$ captures aleatoric uncertainty by computing expected entropy $\mathcal{H}(p(S\mid f,\theta^{*}))$, reflecting noise inherent in imaging or anatomy. For efficiency, the Laplace approximation uses a low-rank Hessian parametrisation. With low-rank matrix $P$, the predictive distribution is $p(\eta\mid f,\theta)=\mathcal{N}(\eta\mid\mu_{\theta},PP^{\top}+D)$, where $D$ is diagonal. This follows methods in [18, 27].
Dropout Segmentation Path. To complement the Laplacian segmentation head, we introduce a Dropout-based segmentation path. The Laplace approximation fits a Gaussian distribution over $\theta_{\text{MAP}}$, producing a Gaussian predictive distribution for segmentation logits and masks. This yields well-calibrated but unimodal uncertainty estimates, reflecting smooth local variations from parameter uncertainty near the mode [10]. However, it may miss multiple plausible segmentation hypotheses or modes, especially with ambiguous or OOD inputs. In contrast, Dropout adds stochasticity by randomly dropping units, effectively sampling from a richer, more diverse space of model configurations, enabling a multimodal predictive distribution over segmentations. Thus, we employ a Dropout Segmentation Head $G(f)$ (Fig. 1), based on MC-dropout [14], which takes embeddings from the Foundation Model encoder as input. From this path, we obtain the logit distribution $\eta_{\text{MC}}$ for sampling logits and a set of uncertainty measures $U_{D,k}$.
Counterfactuals and Uncertainty Calibration. We propose a hybrid uncertainty estimation strategy that fuses outputs from two complementary segmentation heads: Laplacian and Dropout. The Laplacian head produces calibrated, unimodal predictions near the MAP estimate, while the Dropout head captures diverse plausible alternatives as natural counterfactuals. Uncertainty is quantified via intra- and inter-path logit variance, measuring disagreement within and between heads. To emphasise clinically meaningful ambiguity, this disagreement is scaled by total entropy through a Hadamard product, suppressing low-signal noise and highlighting regions where both variance and entropy indicate instability, especially under OOD conditions. Though Dropout can be overconfident in OOD scenarios [21], we leverage this by fusing it with the Laplacian path to enhance OOD sensitivity. The fused uncertainty map is defined as $U_{\text{fused}}=\sqrt{\mathrm{Var}(\eta^{L}-\eta^{D})}\circ\mathcal{H}(\eta^{L}-\eta^{D})\cdot\mathrm{Var}(\tilde{U}_{L/D,k}),$ where $\eta^{L}$ and $\eta^{D}$ are logits from the Laplacian and Dropout paths, respectively, and $\tilde{U}_{L/D,k}$ is an auxiliary pixel-wise uncertainty measure (e.g., mutual information). Here, $\mathrm{Var}(\tilde{U}_{L/D,k})$ captures global pixel-level uncertainty dispersion as a weighting factor. This formulation combines local disagreement, entropy, and global variation, amplifying signals in clinically ambiguous regions while suppressing noise. The fusion yields interpretable uncertainty maps that aid clinical decision-making by highlighting low-confidence areas from epistemic or aleatoric uncertainty, helping sonographers decide when to re-scan, refer, or flag abnormalities in challenging fetal ultrasound cases.